# thanks: Equal contributionthanks: Equal contribution

### v4

@@ I Introduction
Quantum simulation has attracted scientific attention since the early 1980s ignited by Richard Feynman’s vision of the necessity of quantum mechanics in the modeling of natural phenomena [1]. Proposed implementations have included atomic, trapped ion, superconducting and photonic platforms [2, 3, 4, 5, 6, 7]. Here we focus on solid-state optical systems due to their potential for growth into large-scale commercial quantum simulators [8, 9, 10, 11, 12].
@@
Here, we explore how all-photonic quantum simulators based on coupled cavity arrays can benefit from an increased interaction rate established in multi-emitter cavity QED. We expand the Jaynes-Cummings-Hubbard approach to the spectrally disordered Tavis-Cummings-Hubbard model (TCHM) [33] and define conditions for polariton creation utilized in all-photonic quantum simulation, aided by the introduction of new localization metrics inspired by condensed matter approaches. Our model targets applications in technologically mature solid-state platforms and is reflective of the state-of-the-art parameters achieved in silicon carbide and diamond color center hosts. We find system limits that can be guiding for future experiments with polaritons in coupled cavity arrays: 1) polaritonic states are easier to create in systems where emitter/cavity interaction exceeds cavity hopping; 2) polariton creation in an array of lossy cavities can be achieved via integration of an increased number of emitters per cavity even in disordered ensembles; 3) as in single cavities, disordered emitter ensembles in coupled cavity arrays can create polaritons by increasing the number of emitters per cavity to reach the cavity protection condition; and 4) disorder in resonances in a coupled cavity array localizes polaritons if the difference between neighboring cavities exceeds the cavity hopping rate.
@@ II The CCA QED model
$H_{\rm TCHM}=$ $\sum_{n=1}^{N}\Big{\{}\omega_{c,n}a^{\dagger}_{n}a_{n}+\sum_{m=1}^{M_{n}}\Big{[}\omega_{e,n,m}\sigma_{n,m}^{+}\sigma_{n,m}^{-}$ (1)
$+g_{n,m}(a_{n}^{\dagger}\sigma_{n,m}^{-}+\sigma_{n,m}^{+}a_{n})\Big{]}-J_{n,n+1}(a_{n}^{\dagger}a_{n+1}+a_{n+1}^{\dagger}a_{n})\Big{\}},$
@@ II.1 Non-disordered CCA QED model
The origin of these spectral features can be decomposed to the QED and the CCA components. The resonant Tavis-Cummings model of $M$ emitters in $N=1$ cavity has the spectrum of two polaritons and $M-1$ degenerate subradiant states, while a CCA of $N>1,M=0$ cavities has a single spectral band of $N$ photonic states. The spectrum of the resonant TCHM is a product of these components as seen in Figure 1. This model has close analogs to the condensed matter models it aims to simulate.
@@ II.2 Condensed matter analogs of TCHM
A more thorough discussion of these analogs and their limitations is available in the Appendix, but we can begin by looking at the TCHM system with identical emitters ($\omega_{e,n,m}=\omega_{e}$). Much of the derivation of eigenstates precisely parallels the calculation of the band structure of tight binding Hamiltonians commonly studied in condensed matter physics. For example, the case of no emitters ($M_{n}=0$) corresponds to the $d=1$ (Bose) Hubbard Model (HM)[35], and the case with a single emitter in each cavity ($M_{n}=1$) to the Periodic Anderson Model (PAM)[36]: the hopping of photons between cavities is analogous to the conduction electrons which hybridize between different sites, and the photon-emitter coupling maps onto the hopping between the conduction electrons and localized orbitals which, like the emitters, do not have a direct intersite (intercavity) overlap. Thus, in the single excitation sector, and for one emitter per cavity, the models are identical. In practice, the polaritonicity (degree of light-matter hybridization) upon which we focus in the sections to follow, holds lessons for singlet formation in which local and itinerant electrons become tightly intertwined in the PAM and Kondo Lattice Model.
@@ II.3 Experimentally-informed TCHM simulation parameters
The spectral inhomogeneity of emitters in fabricated devices, the main study of our model, has been characterized as $\Delta\sim$ 10 GHz for a variety of emitters in silicon carbide and diamond [38, 39]. We represent this parameter through its relation to the collective coupling rate $g\sqrt{M}$ in a cavity, spanning the spectral inhomogeneity across a range of values. The spectral disorder is implemented by sampling emitter angular frequency, $\omega_{e}$, from a Gaussian distribution $P(\omega_{e})=\frac{1}{\sqrt{2\pi}\sigma}{\rm exp}\{-\frac{(\omega_{e}-\omega_{c})^{2}}{2\sigma^{2}}\}$ centered at $\omega_{c}$ with a width of $\Delta=2\sigma$. It is worth noting that the vibronic resonances are three orders of magnitude larger than the inhomogeneous broadening, for example 8.7 THz for the silicon vacancy in 4H-SiC [40, 41], therefore the phonon side band is not expected to play a part in the collective emitter-cavity coupling process. Emitter lifetime in color centers is usually in the 1-15 ns range [42], we select the value $\gamma/2\pi$ = 1/5.8 GHz as representative. Due to $\gamma$ being the lowest rate in the system, its minimal variations among emitters [39] affect the system only marginally, therefore we assume it has a constant value.
@@ III Effects of disorder on polariton formation
It is in general computationally expensive to solve the Lindbladian master equation to obtain exact simulation results [45]. As such, we are restricted to simulating only very small scale systems ($\sim 6$ elements total) even in the low excitation regime. The results of these exact simulations are available in the Appendix. On the other hand, the effective Hamiltonian ($H_{\rm EFF}$) uses the established non-hermitian effective approach to modeling Hamiltonians that are too resource intensive for the current state of the art classical computers to solve. Its approximation effectiveness is limited to the single-excitation regime, which is suitable for our exploration. Taking the effective Hamiltonian approach a step further, we also introduce the nodal and polaritonic participation ratios ($P_{N}$ and $P_{P}$, respectively). This method is derived from the condensed matter participation ratio metrics [46] used to quantify a system’s localization properties. The $P_{N}$ and $P_{P}$ metrics are applied to eigenstates of the $H_{\rm EFF}$ to quantify the delocalization and light-matter hybridization of the wavefunction, respectively.
@@
$H_{\rm EFF}=H_{\rm TCHM}$ $-\frac{i}{2}\sum_{n=1}^{N}\Big{\{}\kappa_{n}a^{\dagger}_{n}a_{n}+\sum_{m=1}^{M_{n}}\gamma_{n,m}\sigma_{n,m}^{+}\sigma_{n,m}^{-}\Big{\}},$ (3)
@@ III.1 The Participation Ratio Approach: Metrics for characterizing disorder
where $\mathcal{N}_{ph,n}=a_{n}^{\dagger}a_{n}$ and $\mathcal{N}_{e,n}=\sum_{m=1}^{M_{n}}\sigma_{n,m}^{+}\sigma_{n,m}^{-}$ are the usual number operators for each state $v_{p}$ representing cavity excitation and the sum of all emitter excitation in a cavity. Like the classic participation ratio, $P_{N}$ is at a minimum (maximum) when the wavefunction is localized (delocalized). Next, we define the polaritonic participation ratio, or polaritonicity,
@@
which is minimized when the wavefunction has completely cavity-like or completely emitter-like character and is maximized for an equal superposition of cavity- and emitter-like components. Note: a wavefunction can be polaritonic even when the cavity and emitter excitations do not belong to the same node.
@@ IV Discussion
V Appendix: Supplementary Information
V.1 Condensed Matter analogies
The resonant cavity-emitter arrays ($\omega_{c,n}=\omega_{c}$, $\omega_{e,n,m}=\omega_{e}$, $g_{n,m}=g$, $J_{n,n+1}=J$), with periodic boundary conditions have a closed form solution for the eigenvalues and eigenvectors. Much of the derivation precisely parallels the calculation of the band structure of tight binding Hamiltonians. The case of no emitters ($M_{n}=0$) corresponds to the $d=1$ (Bose) Hubbard Model (HM), and the case with a single emitter in each cavity ($M_{n}=1$) to the Periodic Anderson Model (PAM). Here we review those results with an emphasis on the condensed matter analogies.
In the case $M_{n}=0$ one can diagonalize $H_{\rm TCHM}$ by introducing momentum creation and destruction operators,
$a_{k}^{\dagger}=\frac{1}{\sqrt{N}}\,\sum_{l=1}^{N}e^{ikl}a_{l}^{\dagger}\hskip 21.68121pta_{k}=\frac{1}{\sqrt{N}}\,\sum_{l=1}^{N}e^{-ikl}a_{l}$ (A1)
The transformation is canonical. $a_{k}^{\dagger}$ and $a_{k}$ obey the same bosonic commutation relations as the original real space operators. $H_{\rm TCHM}$ is diagonal,
$H_{\rm TCHM}=\sum_{k}\big{(}\omega_{c}-2J\,{\rm cos}\,k\big{)}\,a_{k}^{\dagger}a_{k}$ (A2)
The eigenenergies are,
$E(k)=\omega_{c}-2J\,{\rm cos}\,k\,\,.$ (A3)
The momenta are discrete, $k=k_{p}=2\pi p/N$ with $p=1,2,3,\cdots N$. In the thermodynamic limit the $E(k)$ form a continuous band with a density of states that diverges at the band edges $E(k)=\omega_{c}\pm 2J$.
In this no-emitter limit, Eq. A2 actually provides the solution for any number of excitations. The many-excitation energies are just sums of the single particle $E(k)$ subject to the photon indistinguishability implied by the commutation relations. This solubility of the many excitation system is unique to the no-emitter limit, as discussed further below.
In the case $M_{n}=M=1$ one can again solve for the eigenvalues of $H_{\rm TCHM}$ by going to momentum space, but only in the single excitation sector. The reason is that the photon and emitter operators do not obey a consistent set of commutation relations. Thus, even though it might appear that $H_{\rm TCHM}$ is soluble since it is quadratic in the operators, it is therefore not possible to do the same sort of canonical transformation to diagonalize. It is most straightforward to define a set of single excitation states which form a basis for the space, and then examine the matrix which arises from the application of $H_{\rm TCHM}$ to each one. For simplicity, we focus on the resonant case where $\omega_{c}=\omega_{e}\equiv\omega_{0}$, but the formulae are straightforward to generalize. The energy eigenvalues are
$E(k)=\omega_{0}-J\,{\rm cos}\,k\pm\sqrt{J^{2}\,{\rm cos}^{2}\,k+g^{2}}\,\,.$ (A4)
The (unnormalized) eigenvectors are,
$\Psi_{\pm}(k)=\left(\begin{array}[]{c}g\\[5.69046pt] J{\rm cos}\,k\mp\sqrt{J^{2}\,{\rm cos}^{2}\,k+g^{2}}\\ \end{array}\right)$ (A7)
so that, while all states are polaritons in the sense of mixing photon and emitter components, the relative weights depend on momentum $k$ and band index $\pm$.
This $M_{n}=M=1$ case easily generalizes to larger $M$. There are again two polariton bands, but with an enhanced photon-emitter hybridization $\sqrt{M}g$,
$E(k)=\omega_{0}-J\,{\rm cos}\,k\pm\sqrt{J^{2}\,{\rm cos}^{2}\,k+Mg^{2}}\,\,.$ (A8)
with a similar $g^{2}\rightarrow Mg^{2}$ change to the eigenvectors of Eq. A7. The remaining $M-1$ bands have purely emitter components, and are dispersionless, $E(k)=\omega_{0}$.
Figure A1: The two polariton bands of a cavity-emitter system for representative parameters $\omega_{c}=\omega_{e}\equiv\omega_{0}=4$, cavity hopping $J=1$, and cavity-emitter coupling $g=0.2$. If the number of emitters $M>1$ the gap between the two bands is enhanced from $\Delta=2g$ to $\Delta=2\sqrt{M}g$ and there are $M-1$ additional flat emitter bands at $E_{\alpha}(k)=\omega_{0}$ in the gap between the two polariton bands.
A useful approximate visualization of the eigenvalues and eigenvector weights is provided by drawing the bands as in Figure A1. For $g=0$ the photons have $E(k)=\omega_{0}-2J\,{\rm cos}k$ and the emitters $E(k)=\omega_{0}$. When these two energy levels are hybridized by $g$ there is a level repulsion at their crossing point at $k=\pm\frac{\pi}{2}$ and an energy gap $\Delta=2g$ opens. The relative photon-emitter compositions of the states can be inferred from the degree to which the polariton energy matches one of the initial ($g=0$) photon or emitter bands. Polariton energies which are close to the original flat $E(k)=\omega_{0}$ emitter band are dominantly made up of emitter excitations, while those close to the original dispersing $\omega_{0}-2J\,{\rm cos}k$ photon band are dominantly cavity excitations.
V.2 Exact simulation results
Figure A2: The top plots show the pump operator acting on Cavity 1 and the bottom set show the pump term acting on Cavity 2. Each line within a subplot refers to the transmission spectrum acquired by probing a particular element of the systems, either Cavities or any of the Emitters. Cavity 1 (blue) is coupled to Emitters 1,1 and 1,2 (green and red respectively). Cavity 2 is coupled to Emitters 2,1 and 2,2 (purple and brown respectively). The rows correspond to different cavity-cavity to cavity-emitter coupling ratios, with $J/g=0.1$ (top row) having the least coupling between cavities, and $J/g=10$ (bottom row) having the most. Moving from left to right increases the amount of disorder introduced via emitter emission wavelength detuning from a nominally zero amount ($\epsilon$), to detuning approximately equal to the cavity-emitter coupling ($g$).
Figure A3: Node occupancy of the lowest energy eigenstate of a system with $N=2$ cavities and $M=2$ emitters per cavity with random $\omega_{e}$ sampled from a Gaussian distribution of width $\Delta$. Single random realization for each $\Delta$. $g$, $\kappa$, $\gamma$, and emitter frequencies are the same as those in Figure A2.
We numerically solve the quantum master equation for a coupled 2 cavity system with each cavity also coupling to two emitters each. The results are shown in Figure A2. Along the top row, which shows the smallest inter-cavity coupling simulations, we see the most nodal localization regardless of emitter dispersion, as shown by only having peaks in Cavity 1 and Emitters 1 and 2 when pumping Cavity 1 and only having peaks in Cavity 2 and Emitters 3 and 4 when pumping Cavity 2. We can also see that the small inter-cavity coupling leads to highly polaritonic systems by comparing the peak heights of each cavity to those of its emitters and finding the ratios are close to unity (0.78 - 0.9). The largest inter-cavity coupling, $J/g=10$ is heavily cavity-like as shown by cavity peak to emitter peak ratios of around 50. The middle row, $J/g=1$ is partially polaritonic and partially photonic with peak ratios in the range 0.3-0.6. By introducing nonzero emitter detuning, we are able to access subradiant states for each of the three detuning values. These states are identified by the zero-width peaks found between the polariton peaks. Subradiant states are states that decay much slower than the polariton peaks and a number of proposals exist for their use in quantum information technologies including, light storage [63] and quantum light generation [30].
V.3 Benchmarking of the Effective Hamiltonian approximation and the participation ratio metrics
To benchmark the approximate Effective Hamiltonian approach against the exact Quantum Master Equation solution, we simulate the systems with same parameters in both models. The Figure A3 shows the lowest energy wavefunction calculated with the Effective Hamiltonian approach, corresponding to the excitations of the lowest energy peaks in the Figure A2. The case of the vanishing spectral disorder shows equal contribution of Cavities 1 and 2, which corresponds to the identical looking plots of the Cavity 1 and Cavity 2 excitations in the QME spectra when Cavity 1 and Cavity 2 are pumped, respectively. The asymmetry of the nodal occupations in Figure A3 for non-vanishing disorder (especially for $J/g=0.1$) matches the non-identicality of the Cavity 1/2, as well as Emitter 1.1,1.2 and Emitter 2.1,2.2 spectra in Figure A2. Localization of the wavefunction for an increasing disorder and $J/g\leq 1$ follow the exact solution trends described in the previous section.
Figure A4: Participation ratios of the lowest energy eigenstate for a system with $N=2$ cavities and $M=2$ emitters per cavity with random $\omega_{e}$ sampled from a Gaussian distribution of width $\Delta$. Single random realization for each $\Delta$. $g=5$, $\kappa=10$, $\gamma=1/5.8$. The emitter frequencies are the same as those in Figure A2.
These parallels are closely described by the nodal $P_{N}$ and the polaritonic $P_{P}$ participation ratio shown in Figure A4. The value of $P_{N}$ (orange) is maximized for a fully node-delocalized wavefunction, and reduces as the wavefunction tends to increasingly excite one cavity and its emitters with an increasing disorder. The value of $P_{P}$ (green) is maximized for the wavefunctions that have equal excitation distribution between cavities and the emitters, and reduces with an increasing photonic interaction (high $J/g$ value) as the cavities become predominantly excited.
We conclude that the Effective Hamiltonian and the participation ratio metrics are suitable for studies of the Tavis-Cummings-Hubbard model.
V.4 Effects of spectral disorder on the TCHM energy spectrum
We utilize the Effective Hamiltonian approach to study medium and large TCHM systems and are especially concerned with the influence of the spectral disorder on the polaritonic and localization properties of the model wavefunctions. The Figure A6 shows the energy spectra of an $N=5,M=3$ system for various regimes of cavity QED to photon hopping ratios. For the vanishing disorder, we clearly see the three components of the spectrum illustrated in the Figure 1: two polaritonic bands with $N$ states and a subradiant band with $N(M-1)$ degenerate states. When cavity QED interaction is significant, a band gap opens between the polariton bands and the subradiant states. With an increasing spectral disorder, the band gap closes and the subradiant state degeneracy is lifted.
Figure A5: Effects of an increasing disorder, $\Delta$, on the energy eigenspectrum for 3 different $J/g$ values. Mean of 100 random realizations; error bars are one standard deviation. $N=5$, $M=3$; $g$, $\kappa$, $\gamma$ same as those in Figure A2
Figure A5: Effects of an increasing disorder, $\Delta$, on the energy eigenspectrum for 3 different $J/g$ values. Mean of 100 random realizations; error bars are one standard deviation. $N=5$, $M=3$; $g$, $\kappa$, $\gamma$ same as those in Figure A2
Figure A6: Effects of an increasing disorder, $\Delta$, on the localization and polaritonicity of the eigenstates for 3 different $J/g$ values. For large $J/g$ there remains two polaritonic states located in the middle of the polariton bands. Mean of 100 random realizations; error bars are one standard deviation. $N=5$, $M=3$; $g$, $\kappa$, $\gamma$ same as those in Figure A2
Figure A6: Effects of an increasing disorder, $\Delta$, on the localization and polaritonicity of the eigenstates for 3 different $J/g$ values. For large $J/g$ there remains two polaritonic states located in the middle of the polariton bands. Mean of 100 random realizations; error bars are one standard deviation. $N=5$, $M=3$; $g$, $\kappa$, $\gamma$ same as those in Figure A2
Further analysis is provided by the corresponding $P_{N}$ and $P_{P}$ values shown in Figure A6. Here, we observe that the polaritonic properties of the highly hybridized states in the polariton bands and the emitter-like states in the subradiant band shift significantly only for the high levels of spectral disorder, defined by the typical cavity protection cutoff $\Delta=g\sqrt{M}$. The increasing localization trend (decreasing $P_{N}$) for most polariton band states with an increasing disorder is evident for all sets of parameters. In contrast, the subradiant states gain a cavity component with an increasing disorder and become more hybridized and delocalized.
While the polaritonicity of the lower (as well as the upper) polariton band reduces for most states with an increasing $J/g$ ratio, the middle state of the band remains highly polaritonic. This state, labeled the most polaritonic state (MPS) in our study, shows that even the systems with high hopping ratio can serve as testbeds for polaritonic physics explorations.

### v5

# supplementary_info

@@ Abstract
Keywords: Cavity QED, Tavis-Cummings Hubbard model, Jaynes-Cummings-Hubbard model, polariton simulation, coupled cavity arrays, nanophotonics
@@ I Introduction
Quantum simulation has attracted scientific attention since the early 1980s ignited by Richard Feynman’s vision of the necessity of quantum mechanics in the modeling of natural phenomena [1]. Proposed implementations have included atomic, trapped ion, superconducting and photonic platforms [2, 3, 4, 5, 6, 7, 8]. Here we focus on solid-state optical systems due to their potential for growth into large-scale commercial quantum simulators [9, 10, 11, 12, 13].
@@
Here, we explore how all-photonic quantum simulators based on coupled cavity arrays can benefit from an increased interaction rate established in multi-emitter cavity QED. We expand the Jaynes-Cummings-Hubbard approach to the spectrally disordered Tavis-Cummings-Hubbard model (TCHM) [34] and define conditions for polariton creation utilized in all-photonic quantum simulation, aided by the introduction of new localization metrics inspired by condensed matter approaches. Our model targets applications in technologically mature solid-state platforms and is reflective of the state-of-the-art parameters achieved in silicon carbide and diamond color center hosts. We find system limits that can be guiding for future experiments with polaritons in coupled cavity arrays: 1) polaritonic states are easier to create in systems where emitter/cavity interaction exceeds cavity hopping; 2) polariton creation in an array of lossy cavities can be achieved via integration of an increased number of emitters per cavity even in disordered ensembles; 3) as in single cavities, disordered emitter ensembles in coupled cavity arrays can create polaritons by increasing the number of emitters per cavity to reach the cavity protection condition; and 4) disorder in resonances in a coupled cavity array localizes polaritons if the difference in frequencies between neighboring cavities exceeds the cavity hopping rate.
@@ II The CCA QED model
$H_{\rm TCHM}=$ $\sum_{n=1}^{N}\Big{\{}\omega^{\phantom{{\dagger}}}_{c,n}a^{\dagger}_{n}a^{\phantom{{\dagger}}}_{n}+\sum_{m=1}^{M_{n}}\Big{[}\omega^{\phantom{+}}_{e,n,m}\sigma_{n,m}^{+}\sigma_{n,m}^{-}$ (1)
$+g^{\phantom{{\dagger}}}_{n,m}(a_{n}^{\dagger}\sigma_{n,m}^{-}+\sigma_{n,m}^{+}a^{\phantom{+}}_{n})\Big{]}-J^{\phantom{{\dagger}}}_{n,n+1}(a_{n}^{\dagger}a^{\phantom{{\dagger}}}_{n+1}+a_{n+1}^{\dagger}a^{\phantom{{\dagger}}}_{n})\Big{\}},$
@@ II.1 Non-disordered CCA QED model
Further discussion of the derivation of these equations for the fully resonant case is presented in Section 1 of the Supplementary Information. The origin of these spectral features can be decomposed to the QED and the CCA components. The resonant Tavis-Cummings model of $M$ emitters in $N=1$ cavity has the spectrum of two polaritons and $M-1$ degenerate subradiant states, while a CCA of $N>1,M=0$ cavities has a single spectral band of $N$ photonic states. The spectrum of the resonant TCHM is a product of these components as seen in Figure 1. This model has close analogs to the condensed matter models it aims to simulate.
@@ II.2 Condensed matter analogs of TCHM
A more thorough discussion of these analogs and their limitations is available in the Supplementary Information, but we can begin by looking at the TCHM system with identical emitters ($\omega_{e,n,m}=\omega_{e}$). Much of the derivation of eigenstates precisely parallels the calculation of the band structure of tight binding Hamiltonians commonly studied in condensed matter physics. For example, the case of no emitters ($M_{n}=0$) corresponds to the $d=1$ (Bose) Hubbard Model (HM)[36], and the case with a single emitter in each cavity ($M_{n}=1$) to the Periodic Anderson Model (PAM)[37]: the hopping of photons between cavities is analogous to the conduction electrons which hybridize between different sites, and the photon-emitter coupling maps onto the hopping between the conduction electrons and localized orbitals which, like the emitters, do not have a direct intersite (intercavity) overlap. Thus, in the single excitation sector, and for one emitter per cavity, the models are identical. In practice, the polaritonicity (degree of light-matter hybridization) upon which we focus in the sections to follow, holds lessons for singlet formation in which local and itinerant electrons become tightly intertwined in the PAM and Kondo Lattice Model.
@@ II.3 Experimentally-informed TCHM simulation parameters
The spectral inhomogeneity of emitters in fabricated devices, the main study of our model, has been characterized as $\Delta\sim$ 10 GHz for a variety of emitters in silicon carbide and diamond [39, 40]. We represent this parameter through its relation to the collective coupling rate $g\sqrt{M}$ in a cavity, spanning the spectral inhomogeneity across a range of values. The spectral disorder is implemented by sampling emitter angular frequency, $\omega_{e}$, from a Gaussian distribution $P(\omega_{e})=\frac{1}{\sqrt{2\pi}\sigma}{\rm exp}\{-\frac{(\omega_{e}-\omega_{c})^{2}}{2\sigma^{2}}\}$ centered at $\omega_{c}$ with a width of $\Delta=2\sigma$. It is worth noting that the vibronic resonances are three orders of magnitude larger than the inhomogeneous broadening, for example 8.7 THz for the silicon vacancy in 4H-SiC [41, 42], therefore the phonon side band is not expected to play a part in the collective emitter-cavity coupling process. Emitter lifetime in color centers is usually in the 1-15 ns range [43], we select the value of the spontaneous emission from the color center, $\gamma/2\pi$ = 1/5.8 GHz, as representative. Due to $\gamma$ being the lowest rate in a color center-based cQED system, its minimal variations among emitters of the same species [40] affect the system only marginally, therefore we assume it has a constant value; this is representative of systems like atoms and color centers.
@@ III Effects of disorder on polariton formation
It is in general computationally expensive to solve the Lindbladian master equation to obtain exact simulation results [46]. As such, we are restricted to simulating only very small scale systems ($\sim 6$ elements total, an element being a single cavity or emitter) even in the low excitation regime. The results of these exact simulations are available in Sections 3 and 4 of the Supplementary Information; the matrix form of the Hamiltonian that describes the system simulated is available in Section 2 of the Supplementary Information. On the other hand, the effective Hamiltonian ($H_{\rm EFF}$) uses the established non-hermitian effective approach to modeling Hamiltonians that are too resource intensive for the current state of the art classical computers to solve. Its approximation effectiveness is limited to the single-excitation regime, which is suitable for our exploration. Taking the effective Hamiltonian approach a step further, we also introduce the nodal and polaritonic participation ratios ($P_{N}$ and $P_{P}$, respectively). This method is derived from the condensed matter participation ratio metrics [47] used to quantify a system’s localization properties. The $P_{N}$ and $P_{P}$ metrics are applied to eigenstates of the $H_{\rm EFF}$ to quantify the delocalization and light-matter hybridization of the wavefunction, respectively.
@@
$H_{\rm EFF}=H_{\rm TCHM}$ $-\frac{i}{2}\sum_{n=1}^{N}\Big{\{}\kappa^{\phantom{{\dagger}}}_{n}a^{\dagger}_{n}a^{\phantom{{\dagger}}}_{n}+\sum_{m=1}^{M_{n}}\gamma^{\phantom{+}}_{n,m}\sigma_{n,m}^{+}\sigma_{n,m}^{-}\Big{\}},$ (3)
@@ III.1 The Participation Ratio Approach: Metrics for characterizing disorder
where $\mathcal{N}_{ph,n}=a_{n}^{\dagger}a^{\phantom{{\dagger}}}_{n}$ and $\mathcal{N}_{e,n}=\sum_{m=1}^{M_{n}}\sigma_{n,m}^{+}\sigma_{n,m}^{-}$ are the usual number operators for each state $v_{p}$ representing cavity excitation and the sum of all emitter excitation in a cavity. Like the classic participation ratio, $P_{N}$ is at a minimum (maximum) when the wavefunction is localized (delocalized). Next, we define the polaritonic participation ratio, or polaritonicity,
@@
which is minimized when the wavefunction has completely cavity-like or completely emitter-like character and is maximized for an equal superposition of cavity- and emitter-like components. Here we assume the character is polaritonic when there is any type of light-matter hybridization whether or not it is coming from the same node and therefore note that a wavefunction can be polaritonic even when the cavity and emitter excitations do not belong to the same node.